OpenStudy (amonoconnor):

How would one factor the this: "(3x^2 - 5x + 2) = 0"? Any and all help is greatly appreciated! :)

OpenStudy (e.mccormick):

What did you try?

OpenStudy (zzr0ck3r):

3*2 = 6 we awnt two numbers that multuply to 6 and add to -5, can you think of any?

OpenStudy (amonoconnor):

(-3x + 1)(-x + 2) = 0 That's the best thing I've gotten so far...

OpenStudy (zzr0ck3r):

so 3*2= 6, we want two numbers that multiply to 6, and add to -5 two such numbers are -3, -2 so 3x^2-5x+2 = 3x^2-3x-2x+2 = 3x(x-1)-2(x-1) = (x-1)(3x-2)

OpenStudy (amonoconnor):

Okay... I get that. :)

OpenStudy (amonoconnor):

Sweet, yep this gets me the final answers I was racking my brain to figure out how to get. Thanks a million!